OpenStudy (anonymous):

2+cos(2x)=3cos(x) on the interval [0/2pi] (I know one answer is 0 but there are three more possible answers)

OpenStudy (anonymous):

rewrite \[\cos(2x)\] as \[2\cos^2(x)-1\] and the solve the quadratic

OpenStudy (anonymous):

Is that another formula you have memorized?

OpenStudy (anonymous):

you will get \[2+2\cos^2(x)-1=3\cos(x)\] and frankly (@polpak) i cannot think of a way to do it without rewriting in terms of one function. ??

OpenStudy (anonymous):

No that's the right way. I'm just envious. I was going to have to derive it from Euler's equation.

OpenStudy (anonymous):

well sure, but you would still have to know what to do right? i mean you would still have to think "i have to write this in term of cosine, so how can i do it?"

OpenStudy (anonymous):

Yes indeed.

myininaya (myininaya):

you don't need euler's formula to derive do you remember cos(x+y)=cos(x)cos(y)-sin(x)sin(y)

OpenStudy (anonymous):

truth is that this is a good one to remember because you notice that it does everything in terms of sine

OpenStudy (anonymous):

@myininaya ... you gotta know something to start right?

OpenStudy (anonymous):

But my head is a small hollow plastic cylinder! How can I remember all those trig identities!

myininaya (myininaya):

right

OpenStudy (anonymous):

well i can remember the laws of exponents if nothing else

myininaya (myininaya):

there is indeed a lot of trig identities

OpenStudy (anonymous):

Euler's formula makes them derivable from one piece of information, but it takes a while to work them out each time.

OpenStudy (anonymous):

unless I'm doing a lot of problems that use the same one, then I can write it down ahead of time and re-use it.

OpenStudy (anonymous):

so let us stipulate that we have decided that we need to write everything in terms of cos(x) then how can we do it? we have to get rid of cos(2x) somehow. if i am stuck, and all i remember is that there is some formula, since i am already on the computer i can google it. true if i was in the desert writing in the sand i would have to derive it. assuming i had a stick but then i wouldn't be on openstudy either

myininaya (myininaya):

hint: you will need the quadratic formula (or factoring is an option)

OpenStudy (anonymous):

I'm not against using formula. Sorry if I sound critical.

OpenStudy (anonymous):

I just can't remember them for beans.

OpenStudy (anonymous):

\[2+2\cos^2(x)-1=3\cos(x)\] \[2\cos^2(x)+1=3\cos(x)\] \[2\cos^2(x)-3\cos(x)+1=0\] \[(\cos(x)-1)(2\cos(x)+1)=0\] \[\cos(x)=1,\cos(x)=-\frac{1}{2}\] and then you are done

myininaya (myininaya):

did you factor right?

myininaya (myininaya):

that doesn't look right to me

OpenStudy (anonymous):

me neither unless i am using them. that is what the textbook is for. of course you can derive it from scratch each time. in fact, i looked it up just know!

OpenStudy (anonymous):

@myinaya probably not

myininaya (myininaya):

\[(\cos(x)-1)(2\cos(x)-1)=0\]

OpenStudy (anonymous):

\[(\cos(x)-1)(2\cos(x)-1)=0\]

OpenStudy (anonymous):

lol

myininaya (myininaya):

i win :)

OpenStudy (anonymous):

You both win. ;)

OpenStudy (anonymous):

i meant "just now" because i don't "just know" it

OpenStudy (anonymous):

Oh good. I can feel a bit less inferior then =)

OpenStudy (anonymous):

myininaya knows it because that is how you integrate \[\int\cos^2(x)dx\]

OpenStudy (anonymous):

@polpak... "inferior"??

myininaya (myininaya):

well i would use cos(x+x) to derive the formula for cos(2x)=2cos^2(x)-1

OpenStudy (anonymous):

yes of course. but then you need remember that one too!

OpenStudy (anonymous):

My memory is bad.

OpenStudy (anonymous):

how bout remembering that \[e^{a+b}=e^a\times e^b\]

myininaya (myininaya):

\[\cos(2x)=\cos(x+x)=\cos(x)\cos(x)-\sin(x)\sin(x) \] \[=\cos^2(x)-\sin^2(x)=\cos^2(x)-(1-\cos^2(x)) \] and then also you need to recall \[\sin^2(x)+\cos^2(x)=1\]

myininaya (myininaya):

\[\cos(2x)=\cos^2(x)-(1-\cos^2(x))=2\cos^2(x)-1\]

OpenStudy (anonymous):

truth is you have to start somewhere. and as usual with math the more you know the easier the explanations are

myininaya (myininaya):

the truth is i never learned euler's formula isn't that weird?

myininaya (myininaya):

none of my teachers ever mentioned it

OpenStudy (anonymous):

so if you know addition formula you can use that, with a bunch of algebra if you know "euler's formula' it is just the laws of exponents

OpenStudy (anonymous):

and probably if you know category theory there is a three word explanation

myininaya (myininaya):

it seems to be like a real awesome formula

OpenStudy (anonymous):

Have you taken differential equations? That's the first time it was officially mentioned and used in one of my classes

OpenStudy (anonymous):

and even there it was a bit of hand-waving and 'you'll understand more when you take real analysis'

myininaya (myininaya):

wait maybe i don't remember lol i do remember something about a solution being cos(x)+isin(x)

myininaya (myininaya):

being or having the form: cos(x)+isin(x)

myininaya (myininaya):

\[=e^{ix} ?\]

OpenStudy (anonymous):

There's a good video lecture that uses it in the MIT OCW course for Differential Equations.

OpenStudy (anonymous):

you can do it with only the basics of trig and complex . just the basics. the fact that you can write \[a+bi=r(\cos(\theta)+i\sin(\theta))=re^{i\theta}\]

myininaya (myininaya):

what is r?

OpenStudy (anonymous):

oh wait, i see no that is wrong. you need to know the power series expansion of \[e^x\] \[\sin(x)\] and \[\cos(x)\]

OpenStudy (anonymous):

\[r-|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

*=

myininaya (myininaya):

ok ! r=|a+bi| figure that

OpenStudy (anonymous):

if you want to do if step by step, write down the power series expansion for \[\sin(x)\] and \[\cos(x)\] then geometry tell you that \[a+bi=r(\cos(\theta)+i\sin(\theta))\]

OpenStudy (anonymous):